Find the area of a right triangle whose two sides are 24cm and 7cm and the perimeter is 56cm.​

Grade:12

1 Answers

Abhishek Gupta
3 years ago
Perimeter=56
a+b+c=56
24+7+c=56
c=25
24,7and 25 makes a right angle triangle  where hypotenuse=25
So area of right angled triangle =1/2*base*height
       =1/2*7*24
       =84
So answer is 84